# Verification-Guided Shielding for Deep Reinforcement Learning

### v1

@@ 4 Verification-Guided Shielding
Next, building on the previous findings, we present our novel verification-driven shielding approach. We devised this method on advances from both formal verification as well as shielding and symbolic representation. Our approach incorporates five steps: (1) domain splitting, after which we perform (2) safe-region verification, and (3) clustering. Subsequently, we (4) encode a symbolic representation of the input domain, which finally allows (5) shielding the agent only on potentially unsafe regions.
@@
In the first step, our method employs an off-the-shelf verification algorithm to identify all the regions in which a given DRL agent is more likely to behave correctly. This process mainly relies on the concept of All-DNN-Verification (Marzari et al., 2023). In essence, this includes pruning the input region in search of all regions in which the trained agent is provably safe, with respect to a set of given requirements. More formally, we search for regions in which the negated (safety) property is UNSAT. An exact solution to this problem would provide the complement of the regions where the agent potentially requires a shield.
However, given that this problem is #P-hard (Marzari et al., 2023), the authors proposed $\epsilon$-ProVe, an algorithm that computes an underapproximation of these safe regions. In more detail, $\epsilon$-ProVe divides the input domain into regions, effectively generating a search tree where each node represents a partition of the input domain; these regions are then analyzed with a sampling approach, that provides the estimated probability on whether the region is safe. The algorithm iteratively splits regions into regions, until it cannot find any counterexamples (i.e., UNSAT assignments to the negated property), in which case the region is approximated as safe. Otherwise, the algorithm heuristically decides whether to declare the entire region as unsafe or continue with the splitting procedure. For a detailed description of $\epsilon$-ProVe, and a discussion regarding the probabilistic guarantees provided, we refer the reader to (Marzari et al., 2023).
@@
Currently, we are left with a division of the input domain into regions, with each region approximated as either safe or unsafe. Although $\epsilon$-ProVe typically provides tight results with high confidence, the approximated nature of the approach is not enough to guarantee absolute correctness, which is the subject matter of this work. To address this gap, we complement the approximated regions by formally verifying the regions previously approximated as safe. Toward this end, we employ Marabou, a sound and complete verification tool (Katz et al., 2019b), which is used to formally certify the safety of the agent only in the regions that are previously approximated as safe. In this second, fined-tuned verification procedure, if a counterexample (SAT) is found in a (mistakenly approximated) safe region, we reclassify the region as unsafe. On the other hand, regions that have already been found to violate the property ($\epsilon$-ProVe returned SAT in the first step), are left untouched, as a valid counterexample was already found. A pseudocode describing this procedure can be found in Appendix E.
@@
After these first two steps, we are left with a sound division of the input space to regions in which the agent provably behaves safely and regions in which it might behave unsafely. Next, we would like to apply our synthesized shield solely on these potentially, unsafe regions. However, this presents a new challenge as the set $\mathcal{S}$ of unsafe regions includes, in practice, a large number of compact regions (e.g., we observed an average cardinality of $\approx 60,000$ in Particle World). While this does not compromise the correctness of our strategy, it raises another problem — checking whether the current input belongs to the set $\mathcal{S}$ introduces significant overhead, potentially mitigating the benefits of our approach. To address this limitation, we employ an additional step in which we cluster the set of unsafe regions and reduce their overall cardinality. Specifically, we employ agglomerative clustering (Ackermann et al., 2014), to concatenate unsafe regions and produce an overapproximation of the unsafe regions. As we later demonstrate, this significantly reduces the number of unsafe regions and, consequently, the overhead for checking whether the current state belongs to $\mathcal{S}$. It is important to emphasize that, although the clustering step has the potential to overapproximate safe regions as unsafe (see Fig. 3), it does not compromise the overall soundness of our approach. At most, the shield may be activated more than strictly necessary.
@@
Finally, we are left with a set of (relatively) few approximated unsafe regions. First, we can synthesize a shield that, when activated, guarantees safety w.r.t. the safety property $\psi$. Next, we can couple the shield with the agent, and at each time-step: (i) efficiently identify if the current input belongs to the potentially unsafe regions, and if so, (ii) temporarily activate the synthesized shield and guarantee the safety, as previously described in Sec. 3. In the remaining (provably) safe regions, the shield can remain deactivated while preserving the formal guarantees, as we already verified that any original decision that the agent makes is safe.
Figure 3: An overview of verification-guided shielding. In step (1) we employ $\epsilon$-ProVe to split the input domain into approximated safe (green) and unsafe regions; these can be further validated with a formal verification tool (2a), which complements the set of approximated unsafe regions with the ones formally found as such (2b). To reduce the cardinality of this set we employ a clustering algorithm (3) and further simplify the encoded results by using symbolic representation (4).
@@ 5 Empirical Evaluation
Offline Procedures. Tab. 3 summarizes the time measured for each of the offline steps w.r.t. the Particle World benchmark. The most time-consuming component is the formal verification (step $2$), taking an average of over two hours; this is not surprising, as in this step the verifier is required to solve many NP-complete problems, per each policy. However, we believe this is a reasonable price as (i) this is an offline step that is executed once; and (ii) this is the step that provides the formal guarantees. On the other hand, $\epsilon$-ProVe (i.e., the splitting procedure in step $1$) runs significantly faster but provides only probabilistic assurances. Both stages are complementary and represent a simulated annealing-like optimization (Kirkpatrick et al., 1983): at first, we approximate and reduce the number of regions on which we can rule out correct behavior, and then, we run a more expensive, formal verification procedure that fine-tunes the remaining regions. The column representing the clustering demonstrates a high variance among the different policies, as there can be significant differences in the number of unsafe regions identified in the previous steps. Still, it is worth noting that even in the worst-case measurement, the time required by this procedure is neglectable when compared to the formal verification step. Finally, the table reports the results of the shield synthesis and the formula reduction steps, which were not particularly time-demanding. These results also align with the hypothesis raised in the previous sections, i.e., that the heavy computational cost of shielding is not related to the offline synthesis time, but rather to the cost of invoking the shield online, before each decision.
@@
Online Invocation. Our main results are presented in Tab. 4. Specifically, we compare the overhead introduced by the shield in two cases: the classic fully-activated shielding approach and our verification-guided approach. The first half of the table reports the analysis on the Particle World environment, while the second one reports the results for Mapless Navigation. In general, both benchmarks demonstrate the merits of our approach in reducing significant overhead, confirming the environment-agnostic nature of our methodology. In more detail, per each model and environment, we compare the portion of time in which the shield was activated during execution. This value is computed by normalizing the number of shield invocations by the total number of actions. While in the first column, the (full) shield is trivially always active, we observe a drastic reduction when our approach has been applied, especially in Mapless Navigation. We also note that there is not necessarily a direct correlation between the size of the unsafe regions and the number of interventions, as the interventions were measured with respect to stochastic executions. Finally, we report the average overhead as a relative value compared to an episode with the shield deactivated, i.e., decisions made without invoking any external components. Not surprisingly, our results show a clear correlation between the active time and the overhead introduced, further motivating this work. In the last column, we summarize the time gain provided by our method with respect to invoking the shield at every time-step.
@@
Limitations. Although these results are encouraging, it is important to acknowledge certain limitations of our approach, mainly inherited from the backend shielding and verification techniques. First, our approach requires a valid encoding of the required properties of interest. This, in turn, assumes access to the environment dynamics and the agent’s transition model, as well as the practitioner’s ability to encode the relevant properties in a logic-based form. In addition, when relying on shielding in unsafe regions, it is important to note that although the shield guarantees adherence to the given requirements, it does not necessarily select the optimal action, among the safe ones. Furthermore, there are some limitations due to the backend DNN verification tools. Specifically, our approach relies on Marabou, which is suitable solely for DNNs employing piecewise linear activations in their hidden layers, hence restricting its applicability to more advanced DRL types (e.g., LSTMs, GRUs (Yu et al., 2019), etc.). However, we believe that these limitations can serve as a foundation for future research.
@@ 6 Related Work
Classic shielding approaches (Alshiekh et al., 2018; Pranger et al., 2021a; b) focus on properties expressed in Boolean LTL and are incompatible with systems pertaining to richer data domains. However, more recently, (Wu et al., 2019) presented the concept of (incomplete) shields for linear arithmetic. To address these limitations, the work of (Rodriguez & Sánchez, 2023; 2024a) proves that, under certain conditions, LTL synthesis with data specifications is decidable via abstraction methods, which can be used for reactive synthesis of expressive shields with e.g., numerical information (Rodriguez et al., 2024). Additional techniques, such as runtime enforcement and supervisory control (Cassandras & Lafortune, 1999; Schneider, 2000; Ligatti et al., 2009; Falcone et al., 2012), share similarities with shielding, however, such methods are incompatible with DRL and general reactive systems, but rather, focus solely on checking software invariants.
@@ 7 Conclusion
In this paper, we combine verification and shielding and propose a novel technique that incorporates the advantages of both these formal approaches. Specifically, we demonstrate how to use formal methods to prune the input space and divide it into safe and (potentially) unsafe regions. While in the first case, we can safely employ the original, shield-less model, in the latter we activate the shield online, and override any potential unsafe action. We extensively evaluate our approach in multiple experiments, and demonstrate that it drastically reduces the overhead of shielding, while maintaining guaranteed safety throughout the whole input domain.
Moving forward, our approach can be extended in various axes. Currently, we employ clustering algorithms to approximate unsafe regions, however, we plan to investigate additional strategies that attain more compact descriptions and reduce the use of shielding even further. Additionally, we plan to incorporate our approach also into the DRL training phase to iteratively robustify the agents prior to deployment. Finally, we plan to explore alternative strategies, such as deep ensembles, to further reduce the need for shield interventions. We believe this work is another step towards the reliable use of DRL in safety-critical domains.
@@ Acknowledgments
The work of Amir and Katz was partially funded by the European Union (ERC, VeriDeL, 101112713). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of Amir was further supported by a scholarship from the Clore Israel Foundation. The work of Rodríguez and Sánchez was funded in part by PRODIGY Project (TED2021-132464B-I00) — funded by MCIN/AEI/10.13039/501100011033/ and the European Union NextGenerationEU/PRTR — by the DECO Project (PID2022-138072OB-I00) — funded by MCIN/AEI/10.13039/501100011033 and by the ESF, as well as by a research grant from Nomadic Labs and the Tezos Foundation.

### v2

@@ 4 Verification-Guided Shielding
Next, building on the previous findings, we present our novel verification-driven shielding approach. We devised this method using recent advances in formal verification, shielding, and symbolic representation. Our approach incorporates five steps: (1) domain splitting, after which we perform (2) safe-region verification, and (3) clustering. Subsequently, we (4) encode a symbolic representation of the input domain, which finally allows (5) shielding the agent only on potentially unsafe regions.
@@
In the first step, our method employs an off-the-shelf verification algorithm to identify all the regions of the state space in which a given DRL agent behaves correctly. This process mainly relies on the concept of All-DNN-Verification (Marzari et al., 2023). In essence, this includes pruning the input region in search of all regions in which the trained agent is provably safe, with respect to a set of given requirements. More formally, we search for regions in which the negated (safety) property is UNSAT. An exact solution to this problem would provide the complement of the regions where the agent potentially requires a shield.
However, given that this problem is #P-hard (Marzari et al., 2023), the authors proposed $\epsilon$-ProVe, an algorithm that computes an underapproximation of these safe regions. In more detail, $\epsilon$-ProVe divides the input domain into regions, effectively generating a search tree where each node represents a partition of the input domain; these regions are then analyzed using a sampling approach that provides an estimated probability that the region is safe. The algorithm iteratively splits regions into subregions, until it cannot find any counterexamples (i.e., UNSAT assignments to the negated property), in which case the region is approximated as safe. Otherwise, the algorithm heuristically decides whether to declare the entire region as unsafe or continue with the splitting procedure. For a detailed description of $\epsilon$-ProVe, and a discussion regarding the probabilistic guarantees provided, we refer the reader to (Marzari et al., 2023).
@@
Subsequently, we are left with a division of the input domain into regions, with each region approximated as either safe or unsafe. Although $\epsilon$-ProVe typically provides tight results with high confidence, the approximated nature of the approach is not enough to guarantee absolute correctness, which is the subject matter of this work. To address this gap, we complement the approximated regions by formally verifying the regions previously approximated as safe. Toward this end, we employ Marabou, a sound and complete verification tool (Katz et al., 2019b; Wu et al., 2024), which is used to formally certify the safety of the agent only in the regions that are previously approximated as safe. In this second, fined-tuned verification procedure, if a counterexample (SAT) is found in a (mistakenly approximated) safe region, we reclassify the region as unsafe. On the other hand, regions that have already been found to violate the property ($\epsilon$-ProVe returned SAT in the first step), are left untouched, as a valid counterexample was already found. A pseudocode describing this procedure can be found in Appendix E.
@@
After these first two steps, we are left with a sound division of the input space into regions in which the agent is provably safe and regions in which there is at least one input configuration that causes the agent to behave unsafely. Next, we would like to apply our synthesized shield solely on these potentially unsafe regions. However, this presents a new challenge as the set $\mathcal{S}$ of unsafe regions includes, in practice, a large number of compact regions (e.g., we observed an average cardinality of $\approx 60,000$ in Particle World). While this does not compromise the correctness of our strategy, it raises another problem — checking whether the current input belongs to the set $\mathcal{S}$ introduces significant overhead, potentially mitigating the benefits of our approach. To address this limitation, we employ an additional step in which we cluster the set of unsafe regions and reduce their overall cardinality. Specifically, we employ agglomerative clustering (Ackermann et al., 2014), to concatenate unsafe regions and produce an overapproximation of the unsafe regions. As we later demonstrate, this significantly reduces the number of unsafe regions and, consequently, the overhead for checking whether the current state belongs to $\mathcal{S}$. It is important to emphasize that, although the clustering step has the potential to overapproximate safe regions as unsafe (see Fig. 3), it does not compromise the overall soundness of our approach. At most, the shield may be activated more than strictly necessary.
@@
Finally, we are left with a set of (relatively) few approximated unsafe regions. First, we can synthesize a shield that, when activated, guarantees safety with respect to the safety property $\psi$. Next, we can couple the shield with the agent, and at each time-step: (i) efficiently identify if the current input belongs to the potentially unsafe regions, and if so, (ii) temporarily activate the synthesized shield and guarantee the safety, as previously described in Sec. 3. In the remaining (provably) safe regions, the shield can remain inactive while preserving the formal guarantees, as we already verified that any original decision that the agent makes is safe.
Figure 3: An overview of verification-guided shielding. In step (1) we employ $\epsilon$-ProVe to split the input domain into approximated safe (green) and unsafe regions; (2a) these can be further validated with a formal verification tool, (2b) which complements the set of approximated unsafe regions with the ones formally found as such. (3) To reduce the cardinality of this set we employ a clustering algorithm and (4) further simplify the encoded results by using symbolic representation.
@@ 5 Empirical Evaluation
Offline Procedures. Tab. 3 summarizes the time measured for each of the offline steps with respect to the Particle World benchmark. The most time-consuming component is the formal verification (step $2$), taking an average of over two hours; this is not surprising, as in this step the verifier is required to solve many NP-complete problems, per each policy. However, we believe this is a reasonable price as (i) this is an offline step that is executed once; and (ii) this is the step that provides the formal guarantees. On the other hand, $\epsilon$-ProVe (i.e., the splitting procedure in step $1$) runs significantly faster but provides only probabilistic assurances. Both stages are complementary and represent a simulated annealing-like optimization (Kirkpatrick et al., 1983): at first, we approximate and reduce the number of regions on which we can rule out correct behavior, and then, we run a more expensive, formal verification procedure that fine-tunes the remaining regions. The column representing the clustering demonstrates a high variance among the different policies, as there can be significant differences in the number of unsafe regions identified in the previous steps. Still, it is worth noting that even in the worst-case measurement, the time required by this procedure is negligible when compared to the formal verification step. Finally, the table reports the results of the shield synthesis and the formula reduction steps, which were not particularly time-demanding (i.e., step 4). These results also align with the hypothesis raised in the previous sections, i.e., that the heavy computational cost of shielding is not related to the offline synthesis time, but rather to the cost of invoking the shield online, before each decision.
@@
Online Invocation. Our main results are presented in Tab. 4. Specifically, we compare the overhead introduced by the shield in two cases: the classic fully-activated shielding approach and our verification-guided approach. The first half of the table reports the analysis on the Particle World environment, while the second one reports the results for Mapless Navigation. In general, both benchmarks demonstrate the merits of our approach in reducing significant overhead, confirming the general environment nature of our methodology. In more detail, per each environment and seed, we compare the portion of time in which the shield was activated during execution. This value is computed by normalizing the number of shield invocations by the total number of actions. While in the first column, the (full) shield is trivially always active, we observe a drastic reduction when our approach has been applied, especially in some Mapless Navigation seeds. We also note that there is not necessarily a direct correlation between the size of the unsafe regions and the number of interventions, as the interventions were measured with respect to stochastic executions. Finally, we report the average computational time overhead as a relative value compared to an episode with the shield deactivated, i.e., decisions made without invoking any external components. Not surprisingly, our results show a clear correlation between the active time and the overhead introduced, further motivating this work. In the last column, we summarize the time gain provided by our method with respect to invoking the shield at every time-step. Note that the gain is not always proportional to the active time; the resulting value also depends on the actual number of steps required to complete a single episode, i.e., the absolute number of calls to the shield.
@@
Limitations. Although these results are encouraging, it is important to acknowledge certain limitations of our approach, mainly inherited from the backend shielding and verification techniques. First, our approach requires a valid encoding of the required properties of interest. This, in turn, assumes access to the environment dynamics and the agent’s transition model, as well as the practitioner’s ability to encode the relevant properties in a logic-based form. In addition, when relying on shielding in unsafe regions, it is important to note that although the shield guarantees adherence to the given requirements, it does not necessarily select the optimal action, among the safe ones. Furthermore, there are some limitations due to the backend DNN verification tools. Specifically, our approach relies on Marabou, which affords only limited support to some activation functions, hence restricting its applicability to some advanced DNN architectures. However, we believe that these limitations can serve as a foundation for future research.
@@ 6 Related Work
Classic shielding approaches (Alshiekh et al., 2018; Pranger et al., 2021a; b) focus on properties expressed in Boolean LTL and are incompatible with systems pertaining to richer data domains. However, more recently, Wu et al. (2019) presented the concept of (incomplete) shields for linear arithmetic. To address these limitations, the work of Rodriguez & Sánchez (2023; 2024a) proves that, under certain conditions, LTL synthesis with data specifications is decidable via abstraction methods, which can be used for reactive synthesis of expressive shields with e.g., numerical information (Rodriguez et al., 2024). Additional techniques, such as runtime enforcement and supervisory control (Cassandras & Lafortune, 1999; Schneider, 2000; Ligatti et al., 2009; Falcone et al., 2012), share similarities with shielding, however, such methods are incompatible with DRL and general reactive systems, but rather, focus solely on checking software invariants.
@@ 7 Conclusion
In this paper, we combine verification and shielding and propose a novel technique that leverages the advantages of both these formal approaches. Specifically, we demonstrate how to use formal methods to prune the input space and divide it into safe and (potentially) unsafe regions. While in the first case, we can safely employ the original, shield-less model, in the latter we activate the shield online, and override any potential unsafe action. We extensively evaluate our approach in multiple experiments, and demonstrate that it drastically reduces the overhead of shielding, while maintaining guaranteed safety throughout the whole input domain.
Moving forward, our approach can be extended along various axes. Currently, we employ clustering algorithms to approximate unsafe regions, however, we plan to investigate additional strategies that attain more compact descriptions and reduce the use of shielding even further. Additionally, we plan to incorporate our approach also into the DRL training phase to iteratively robustify the agents prior to deployment. Finally, we plan to explore alternative strategies, such as deep ensembles, to further reduce the need for shield interventions. We believe this work is another step towards the reliable use of DRL in safety-critical domains.
@@ Acknowledgments
The work of Corsi and Fox was funded in part by the National Science Foundation (Award #2321786). The work of Amir and Katz was partially funded by the European Union (ERC, VeriDeL, 101112713). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. The work of Amir was further supported by a scholarship from the Clore Israel Foundation. The work of Rodríguez and Sánchez was funded in part by PRODIGY Project (TED2021-132464B-I00) — funded by MCIN/AEI/10.13039/501100011033/ and the European Union NextGenerationEU/PRTR — by the DECO Project (PID2022-138072OB-I00) — funded by MCIN/AEI/10.13039/501100011033 and by the ESF, as well as by a research grant from Nomadic Labs and the Tezos Foundation.